OpenStudy (anonymous):

y=x+1;A(1,0) is this point on the line?

OpenStudy (anonymous):

Can anyone tell me what A stands for? Sorry this is only my second day of geometry.

jimthompson5910 (jim_thompson5910):

A just a point, or the name of the point

jimthompson5910 (jim_thompson5910):

You can call it anything you want, the name A has no meaning or significance

jimthompson5910 (jim_thompson5910):

If you want, you can erase the "A" to get (1,0) and you'd still have the same question.

jimthompson5910 (jim_thompson5910):

So is the point (1,0) on the line y = x+1?

OpenStudy (anonymous):

I guess?? I think thats what it's asking

jimthompson5910 (jim_thompson5910):

y = x+1 0 = 1+1 ... Plug in (1,0), basically x = 1 and y = 0 0 = 2 Since that equation is false, (1,0) is NOT a solution So (1,0) is NOT on the line

OpenStudy (anonymous):

So you plugged in 1 for y and 0 for x?

jimthompson5910 (jim_thompson5910):

1 for x

jimthompson5910 (jim_thompson5910):

0 for y

OpenStudy (anonymous):

So the first number in the parenthesis will always be x and the second will be y?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

(x,y) is any point

OpenStudy (anonymous):

okay thanks soooo much just to be certain after you plugged everything in it would be like 0=2 and thats why it's no?

jimthompson5910 (jim_thompson5910):

yes that's exactly why

OpenStudy (anonymous):

thaks so much do you get on this website often?

jimthompson5910 (jim_thompson5910):

i do

OpenStudy (anonymous):

okay because I'm sure I'll be needing more help in the future

jimthompson5910 (jim_thompson5910):

alright, I look forward to helping you in the future if you need it

OpenStudy (anonymous):

thanks again

jimthompson5910 (jim_thompson5910):

np